# Middo: Model-Informed Dynamic Data Optimization for Enhanced LLM Fine-Tuning via Closed-Loop Learning

### v1

@@ Abstract
Supervised Fine-Tuning (SFT) Large Language Models (LLM) fundamentally rely on high-quality training data. While data selection and data synthesis are two common strategies to improve data quality, existing approaches often face limitations in static dataset curation that fail to adapt to evolving model capabilities. In this paper, we introduce Middo, a self-evolving Model-informed dynamic data optimization framework that uses model-aware data selection and context-preserving data refinement. Unlike conventional one-off filtering/synthesis methods, our framework establishes a closed-loop optimization system: (1) A self-referential diagnostic module proactively identifies suboptimal samples through tri-axial model signals - loss patterns (complexity), embedding cluster dynamics (diversity), and self-alignment scores (quality); (2) An adaptive optimization engine then transforms suboptimal samples into pedagogically valuable training points while preserving semantic integrity; (3) This optimization process continuously evolves with model capability through dynamic learning principles. Experiments on multiple benchmarks demonstrate that our Middo consistently enhances the quality of seed data and boosts LLM’s performance with improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. This work establishes a new paradigm for sustainable LLM training through dynamic human-AI co-evolution of data and models.
Middo: Model-Informed Dynamic Data Optimization for Enhanced LLM Fine-Tuning via Closed-Loop Learning
Zinan Tang1,2, Xin Gao2, Qizhi Pei2,3, Zhuoshi Pan2,4 Mengzhang Cai2, Jiang Wu2, Conghui He2, Lijun Wu2thanks:  Corresponding authors. 1School of Computer Science, Beijing University of Posts and Telecommunications 2Shanghai Artificial Intelligence Laboratory  3Renmin University of China  4Tsinghua University tangzinan@bupt.edu.cn{heconghui,wulijun}@pjlab.org.cn
@@
1 Introduction
Large Language Models (LLMs) have revolutionized artificial intelligence by achieving state-of-the-art performance across diverse domains, from natural language understanding (Zhou et al., 2023c; Hendrycks et al., 2021a) to mathematical reasoning (Cobbe et al., 2021; Hendrycks et al., 2021b) and code generation (Chen et al., 2021; Austin et al., 2021). This success is largely attributed to Supervised Fine-Tuning (SFT), where models undergo rigorous training on high-quality, human-aligned datasets to ensure outputs closely match human expectations. Crucially, the quality of these datasets directly dictates the model’s ultimate capabilities: noisy or suboptimal training data can lead to degraded performance, while meticulously curated data unlocks advanced reasoning, generalization, and robustness. As LLMs scale, the adage “garbage in, garbage out” becomes increasingly important—highlighting the urgent need for systematic methods to optimize training data quality.
Existing approaches primarily fall into two categories to improve data quality: data selection (Cao et al., 2024; Zhou et al., 2023b; Li et al., 2024d; Jia et al., 2024; Zhou et al., 2024; Li et al., 2024e, a) and data synthesis (Dai et al., 2023; WANG et al., 2023; Mukherjee et al., 2023; Xu et al., 2025; Liu et al., 2024a). Data selection methods filter raw datasets using heuristic rules (e.g., length filters) (Zhao et al., 2024a) or statistical metrics like perplexity (PPL) (Liu et al., 2024a) and Instruction-Following Difficulty (IFD) (Li et al., 2024c) to retain “high-quality” samples. Conversely, data synthesis leverages advanced LLMs (e.g., GPT-4 (Achiam et al., 2023)) to generate new training examples, often through prompting or distillation (Li et al., 2024f). While both strategies improve data quality, they suffer from critical limitations. Selection methods are typically static, applying fixed criteria that ignore the evolving needs of the model during training. Similarly, synthesis approaches often discard original data, wasting potentially valuable information, and risk generating distributional narrow or redundant examples. These one-time data curation methods fail to adaptively refine data along with the model’s progress.
@@ 1 Introduction
Experiments across multiple benchmarks demonstrate Middo ’s effectiveness especially on low-quality datasets. Models trained with Middo optimized data achieve consistent performance gains over baselines, improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. Notably, Middo trained models exhibit stronger abilities to addresses hard problems, solving more than three times the number of challenging test problems (e.g., MATH, GPQA) compared to models trained on static datasets. These results validate that sustainable LLM advancement requires co-evolving data and models—a paradigm shift from today’s disjointed curation practices.
@@ 2.1 Synthetic Data Generation
Synthetic data generation is a key technique for augmenting LLM fine-tuning. Early methods (Edunov et al., 2018; Wieting and Gimpel, 2018) introduce perturbation-based approaches to enhance data diversity, using character-level (Belinkov and Bisk, 2018) and word-level (Wei and Zou, 2019) modifications. These methods rely on fixed transformation rules, limiting adaptability.
LLMs have been leveraged for scalable data synthesis (Sudalairaj et al., 2024; Jung et al., 2024; Dai et al., 2023; WANG et al., 2023; Mukherjee et al., 2023; Xu et al., 2025; Liu et al., 2024a; Li et al., 2025). Self-instruct methods (Wang et al., 2023) generate instruction-response pairs, while Evol-Instruct (Xu et al., 2024) and Auto-Evol-Instruct (Zeng et al., 2024) refine data complexity iteratively. However, these methods remain static, failing to adapt as models improve. Recent approaches integrate model feedback into data generation (Anonymous, 2025a; Cao et al., 2025; Anonymous, 2025b; Li et al., 2024f), incorporating student model signals for adaptive synthesis. LLM2LLM (Lee et al., 2024) is an iterative data augmentation strategy that enhances low-data fine-tuning by using a teacher LLM to generate synthetic training data from incorrect student LLM predictions and I-SHEEP (Park et al., 2024) uses iterative self-enhancement paradigm.
@@ 2.2 Data Selection
Data selection is crucial for LLM fine-tuning, as high-quality and informative data directly impacts model performance (Zhou et al., 2023a; Xu et al., 2023). Early heuristic-based methods rely on surface-level statistics like item frequency (Raffel et al., 2020) and repetition count (Laurençon et al., 2022), but lack adaptability to model evolution.
Recent work explores LLM-driven data selection, optimizing for quality, diversity, and complexity (Cao et al., 2024; Zhou et al., 2023b; Li et al., 2024d; Jia et al., 2024; Zhou et al., 2024; Li et al., 2024e, a; Du et al., 2023; Kung et al., 2023). Instruction-Following Difficulty (IFD) metric (Li et al., 2024c) enables models to self-select training instances, while other methods (Yu et al., 2024; Colombo et al., 2024; Lu et al., 2024; Zhao et al., 2024b) use LLM self-assessment for efficiency. Further advancements integrate LLM-based evaluation mechanisms. AlpaGasus (Chen et al., 2024) and LIFT (Xu et al., 2023) use structured prompts for data assessment, while DEITA (Liu et al., 2024b) introduces a multi-dimensional scoring system based on complexity and quality.
@@ 3.1 Middo Pipeline
At each iteration, these parallel signal analyzers jointly select suboptimal samples, which are then regenerated through context-aware synthesis—preserving original semantic intent while enhancing pedagogical value. The refined dataset immediately feeds back into model training, creating a dynamic feedback loop where improved model capabilities inform subsequent optimization cycles. Notably, the optimized dataset remains similar data size, without extending large data synthesis, leading an efficient data optimization. This self-referential mechanism ensures continuous alignment between data characteristics and model evolution. Following sections systematically elaborate on the implementation of each signal-specific optimization module and their synergistic integration.
@@ 3.2 Loss Patterns: Complexity Optimization
Complexity reflects the “difficulty” or “compositionality” of data. A good dataset usually requires smooth complexity distribution of data for training. Therefore, we introduce Loss Patterns, which targets overly challenging samples by modifying them to maintain a balanced and learnable training set (Zhao et al., 2024b). During fine-tuning, the loss for a sample $(X_{i},Y_{i})$ is computed as the likelihood of predicting successive tokens given the instruction $X_{i}$ and its context. We denote the loss before and after training as $\mathcal{L}_{\text{pre}}(X_{i},Y_{i})$ and $\mathcal{L}_{\text{post}}(X_{i},Y_{i})$, respectively.
Intuitively, we consider both the loss before and after training to select the complex data. Specifically, we classify samples based on their loss evolution: samples with both low $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ are considered easy, while those with high values in both metrics remain difficult, indicating excessive complexity. A sample is included in the complex subset $\mathcal{D}^{\text{hard}}$ if its $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ both exceed the thresholds $\tau_{\text{pre}}$ and $\tau_{\text{post}}$, respectively. For adaptive refinement, the thresholds are dynamically computed. See Appendix B.3 for details on the dynamic threshold settings used throughout the paper.
@@
For complex data optimization, instead of discarding difficult samples, we transform them to be simpler with more manageable forms. Specifically, we replace samples in $\mathcal{D}^{\text{hard}}$ with their simplified counterparts, $\mathcal{D}^{\text{hard}^{\prime}}$. This is achieved by an automatic process in which a LLM analyzes and summarizes the complex instructions (Zeng et al., 2024), then simplifies them step by step while preserving the core educational content. An example is shown in Appendix Figure 9. This iterative transformation process updates the dataset by replacing overly complex samples with refined versions that offer more effective training samples. As training continues, this adaptive approach ensures a continuous alignment between data complexity and model capability.
@@ 3.3 Embedding Cluster Dynamics: Diversity Optimization
Diversity is crucial for ensuring broad concept coverage and a uniform data distribution. Embedding Cluster Dynamics identifies sparse data points that signal underrepresented regions in the dataset. We use the embedding cluster dynamics to select data as follows: We extract sentence embeddings from the last hidden layer $\mathcal{H}^{(L)}$ of the model trained in the previous iteration, using average pooling, then compute the cosine similarity between each data point and find the $k$-nearest neighbors $\mathcal{N}_{k}(X_{i})$ for each data $X_{i}$. A lower average cosine similarity among these neighbors $\mathcal{N}_{k}(X_{i})$ indicates the data is positioned in a sparser region. Thus, the data points whose average cosine similarity score (diversity score) below a threshold are selected for optimization.
@@ 3.4 Self-alignment Scores: Quality Optimization
High-quality data is essential for fine-tuning, as poor-quality samples can degrade performance (Zhou et al., 2023a). To reduce manual annotation costs, many approaches use the LLM-as-a-Judge paradigm (Chen et al., 2024; Xu et al., 2023). We enhance this by leveraging the fine-tuned model itself to assess data quality via Self-alignment Scores. Specifically, for each instruction-response pair $(X_{i},Y_{i})$ in the dataset $\mathcal{D}$, the model computes scores $\mathcal{S}{\pi}^{\text{ins.}}(X_{i})$ for instruction evaluation and $\mathcal{S}{\pi}^{\text{res.}}(X_{i},Y_{i})$ for instruction-response pair evaluation based on three key metrics $\pi$ from AlignBench (Liu et al., 2024c): Clarity, Completeness, and Factuality. The final quality score $\mathcal{S}(X_{i},Y_{i})$ is obtained by averaging these individual scores. These samples with scores below a similar dynamic threshold are identified as low-quality, forming the seed dataset $\mathcal{D}^{\text{low}}$.
@@
In each iteration, after the above three data selection and refinement progress, the optimized dataset is then feed back for next round model training.
@@ 4 Experiment
4.1 Settings
Data Optimization Configurations.
We conduct optimization on the Alpaca (Taori et al., 2023) and WizardLM (Xu et al., 2024) datasets. To ensure a fair comparison, we also completely rewrite the responses of Alpaca using GPT-4o-mini as seed data for the experiments. For Embedding Cluster Dynamics, we set the number of neighbors to $k=2$. Each dataset undergoes three iterations of optimization. A detailed analysis of the computational cost of Middo is provided in Appendix A, and the effects of the number of neighbors and iteration counts are discussed in Appendix B. We generate data by GPT-4o-mini, setting both temperature and top_p to 1.0 to ensure diversity.
Training and Evaluation Settings.
We fine-tune LLaMA-3.1-8B (Dubey et al., 2024) and Mistral-7B-v0.3 (Jiang et al., 2023) using LLaMA-Factory (Zheng et al., 2024). For each iteration of Middo’s optimization, the base model is fine-tuned for one epoch using full-parameter SFT solely on the dataset refined in that specific iteration. By retraining a model with each newly optimized dataset, rather than continuing training from the model state of a previous iteration, we mitigate the risk of overfitting to the evolving data (Lee et al., 2024; Anonymous, 2025b). Evaluation is conducted using OpenCompass (Contributors, 2023), with vLLM (Kwon et al., 2023) for batch inference acceleration. All hyperparameters are detailed in Appendix C.5. We assess general knowledge on IFEval (Zhou et al., 2023c) and MMLU (Hendrycks et al., 2021a), mathematical problem-solving on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), code generation on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), and commonsense reasoning on Hellaswag (Zellers et al., 2019) and GPQA (Rein et al., 2024).
4.2 Main Results
@@
Alpaca-4o-mini Middo Only 8.8kk 43.47 40.78 65.20 15.58 51.83 47.60 58.65 17.68 42.60
Middo 63.1kk 44.69 47.96 57.62 18.50 52.44 45.40 57.37 19.70 42.96
Table 2: Results of Middo compared to other baseline methods. The best and second best results are highlighted in bold and underlined, respectively. The size of the dataset is the number of examples used for training. Our method outperforms all baselines in the average score. *Note that 0.00 indicates that the method did not solve any examples.
Dynamic Iterative Improvement.
Further Validation on 4o-mini Rewritten Data.
Steady improvements observed on the 4o-mini rewritten Alpaca dataset—averaging a 2.2% increase overall, with MMLU showing an impressive 11.87% boost—demonstrate that these gains are not merely a result of using 4o-mini data. This illustrates that our framework intrinsically enhances data quality and model performance. Importantly, we achieve these improvements without resorting to stronger variants such as GPT-4o (Hurst et al., 2024), reinforcing the robustness and general applicability of our method.
Initial Dataset Quality.
Our experiments reveal that higher-quality datasets require fewer modifications to reach optimal performance. For instance, while the Alpaca dataset achieve peak performance at third iterations, the 4o-mini rewritten Alpaca required only two iterations, and the Wizard dataset reaches its best performance in just one round.
Comparison with Other Works.
We compare Middo with both existing data selection (Alpaca-clean(Ruebsamen, 2023), Superfiltering (Li et al., 2024b), Long (Zhao et al., 2024a), AlpaGasus (Chen et al., 2024)) and data augmentation (Alpaca-GPT4 (Peng et al., 2023), I-SHEEP (Anonymous, 2025b), WizardLM (Xu et al., 2024)) methods on the Alpaca dataset.
We use the optimal dataset obtained through Middo from Alpaca for comparison with other baselines. Additionally, to ensure a relatively fair comparison with data selection methods, we include a dataset that only uses the optimized data without incorporating any unoptimized samples, referred as “Middo Optimize Only”, to isolate the effect of the optimization process and make a direct comparison with data selection approaches.
Results in Table 2 shows our method achieves the highest average score of $42.96$, outperforming all other approaches. Notably, even when using only the optimized subset Middo Optimize Only, our method delivers a robust average score of $42.6$. This demonstrates that data size is not the main factor influencing iterative improvement, but are inherent to the effectiveness of our dynamic data selection and optimization process.

### v2

# 1]OpenDataLab 2]Shanghai Artificial Intelligence Laboratory

@@ Abstract
Supervised Fine-Tuning (SFT) Large Language Models (LLMs) fundamentally rely on high-quality training data. While data selection and data synthesis are two common strategies to improve data quality, existing approaches often face limitations in static dataset curation that fail to adapt to evolving model capabilities. In this paper, we introduce Middo, a self-evolving Model-informed dynamic data optimization framework that uses model-aware data selection and context-preserving data refinement. Unlike conventional one-off filtering/synthesis methods, our framework establishes a closed-loop optimization system: (1) A self-referential diagnostic module proactively identifies suboptimal samples through tri-axial model signals-loss patterns (complexity), embedding cluster dynamics (diversity), and self-alignment scores (quality); (2) An adaptive optimization engine then transforms suboptimal samples into pedagogically valuable training points while preserving semantic integrity; (3) This optimization process continuously evolves with the model’s capability through dynamic learning principles. Experiments on multiple benchmarks demonstrate that our Middo consistently enhances the quality of seed data and boosts LLMs’ performance, improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. This work establishes a new paradigm for sustainable LLM training through dynamic human-AI co-evolution of data and models.
[…]
1 Introduction
Large Language Models (LLMs) have revolutionized artificial intelligence by achieving state-of-the-art performance across diverse domains, from natural language understanding [67, 17] to mathematical reasoning [7, 16] and code generation [6, 2]. This success is largely attributed to Supervised Fine-Tuning (SFT), where models undergo rigorous training on high-quality, human-aligned datasets to ensure outputs closely match human expectations. Crucially, the quality of these datasets directly dictates the model’s ultimate capabilities: noisy or suboptimal training data can lead to degraded performance, while meticulously curated data unlocks advanced reasoning, generalization, and robustness. As LLMs scale, the adage “garbage in, garbage out” becomes increasingly important—highlighting the urgent need for systematic methods to optimize training data quality.
Existing approaches primarily fall into two categories to improve data quality: data selection [4, 65, 31, 19, 66, 32, 28] and data synthesis [10, 51, 40, 57, 35, 14]. Data selection methods filter raw datasets using heuristic rules (e.g., length filters) [61] or statistical metrics like perplexity (PPL) [35] and Instruction-Following Difficulty (IFD) [30] to retain “high-quality” samples. Conversely, data synthesis leverages advanced LLMs (e.g., GPT-4 [1]) to generate new training examples, often through prompting or distillation [33]. While both strategies improve data quality, they suffer from critical limitations. Selection methods are typically static, applying fixed criteria that ignore the evolving needs of the model during training. Similarly, synthesis approaches often discard original data, wasting potentially valuable information, and risk generating distributionally narrow or redundant examples. These one-time data curation methods fail to adaptively refine data along with the model’s progress.
@@ 1 Introduction
Experiments across multiple benchmarks demonstrate Middo ’s effectiveness especially on low-quality datasets. Models trained with Middo optimized data achieve consistent performance gains over baselines, improving accuracy by $7.15\%$ on average while maintaining the original dataset scale. Notably, Middo-trained models exhibit stronger abilities to address hard problems, solving more than three times the number of challenging test problems (e.g., MATH, GPQA) compared to models trained on static datasets. These results validate that sustainable LLM advancement requires co-evolving data and models—a paradigm shift from today’s disjointed curation practices.
@@ 2.1 Synthetic Data Generation
Synthetic data generation is a key technique for augmenting LLM fine-tuning. Early methods [13, 54] introduce perturbation-based approaches to enhance data diversity, using character-level [3] and word-level [53] modifications. These methods rely on fixed transformation rules, limiting adaptability.
LLMs have been leveraged for scalable data synthesis [48, 21, 10, 51, 40, 57, 35, 27]. Self-instruct methods [52] generate instruction-response pairs, while Evol-Instruct [55] and Auto-Evol-Instruct [60] refine data complexity iteratively. However, these methods remain static, failing to adapt as models improve. Recent approaches integrate model feedback into data generation [22, 39, 34, 33], incorporating student model signals for adaptive synthesis. LLM2LLM [26] is an iterative data augmentation strategy that enhances low-data fine-tuning by using a teacher LLM to generate synthetic training data from incorrect student LLM predictions and I-SHEEP [41] uses an iterative self-enhancement paradigm.
@@ 2.2 Data Selection
Data selection is crucial for LLM fine-tuning, as high-quality and informative data directly impacts model performance [64, 56]. Early heuristic-based methods rely on surface-level statistics like item frequency [44] and repetition count [25], but they also lack adaptability to model evolution.
Recent work explores LLM-driven data selection, optimizing for quality, diversity, and complexity [4, 65, 31, 19, 66, 32, 28, 11, 23]. The IFD metric [30] enables models to self-select training instances by comparing loss with and without the instruction, while other methods [58, 8, 38, 62] use LLM self-assessment for efficiency. Further advancements integrate LLM-based evaluation mechanisms. AlpaGasus [5] and LIFT [56] use structured prompts for data assessment, while DEITA [36] introduces a multi-dimensional scoring system based on complexity and quality.
@@ 3.1 Middo Pipeline
At each iteration, these parallel signal analyzers jointly select suboptimal samples, which are then regenerated through context-aware synthesis—preserving original semantic intent while enhancing pedagogical value. The refined dataset immediately feeds back into model training, creating a dynamic feedback loop where improved model capabilities inform subsequent optimization cycles. Notably, the optimized dataset remains similar in data size, without extending large data synthesis, leading to an efficient data optimization. This self-referential mechanism ensures continuous alignment between data characteristics and model evolution. The following sections systematically elaborate on the implementation of each signal-specific optimization module and their synergistic integration.
@@ 3.2 Loss Patterns: Complexity Optimization
Complexity reflects the “difficulty” or “compositionality” of data. A good dataset usually requires a smooth complexity distribution of data for training [15, 47]. Therefore, we introduce Loss Patterns, which targets overly challenging samples by modifying them to maintain a balanced and learnable training set [62]. During fine-tuning, the loss for a sample $(X_{i},Y_{i})$ is computed as the likelihood of predicting successive tokens given the instruction $X_{i}$ and its context. We denote the loss before and after training as $\mathcal{L}_{\text{pre}}(X_{i},Y_{i})$ and $\mathcal{L}_{\text{post}}(X_{i},Y_{i})$, respectively.
Intuitively, we consider both the loss before and after training to select the complex data. Specifically, we classify samples based on their loss evolution: samples with both low $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ are considered easy, while those with high values in both metrics remain difficult, indicating excessive complexity. A sample is included in the complex subset $\mathcal{D}^{\text{hard}}$ if its $\mathcal{L}_{\text{pre}}$ and $\mathcal{L}_{\text{post}}$ both exceed the thresholds $\tau_{\text{pre}}$ and $\tau_{\text{post}}$, respectively. For adaptive refinement, the thresholds are dynamically computed. See Appendix 8.3 for details on the dynamic threshold settings used throughout the paper.
@@
For complex data optimization, instead of discarding difficult samples, we transform them into simpler, more manageable forms. Specifically, we replace samples in $\mathcal{D}^{\text{hard}}$ with their simplified counterparts, $\mathcal{D}^{\text{hard}^{\prime}}$. This is achieved by an automatic process in which a LLM analyzes and summarizes the complex instructions [60], then simplifies them step by step while preserving the core educational content. An example is shown in Appendix Figure 11. This iterative transformation process updates the dataset by replacing overly complex samples with refined versions that offer more effective training samples. As training continues, this adaptive approach ensures a continuous alignment between data complexity and model capability.
@@ 3.3 Embedding Cluster Dynamics: Diversity Optimization
Diversity is crucial for ensuring broad concept coverage and a uniform data distribution. Embedding Cluster Dynamics identifies sparse data points that signal underrepresented regions in the dataset. We extract sentence embeddings from the last hidden layer $\mathcal{H}^{(L)}$ of the model trained in the previous iteration, using average pooling, then compute the cosine similarity between each data point and find the $k$-nearest neighbors $\mathcal{N}_{k}(X_{i})$ for each data $X_{i}$. A lower average cosine similarity among these neighbors $\mathcal{N}_{k}(X_{i})$ indicates the data is positioned in a sparser region. Thus, the data points whose average cosine similarity score (diversity score) is below a threshold are selected for optimization.
@@ 3.4 Self-alignment Scores: Quality Optimization
High-quality data is essential for fine-tuning, as poor-quality samples can degrade performance [64]. To reduce manual annotation costs, many approaches use the LLM-as-a-Judge paradigm [5, 56]. To achieve this, instead of relying on an external judge, we leverage the fine-tuned model itself to assess data quality via Self-alignment Scores, effectively incorporating the model’s own feedback. Specifically, for each instruction-response pair $(X_{i},Y_{i})$ in $\mathcal{D}$, the model generate scores $\mathcal{S}{\pi}^{\text{ins.}}(X_{i})$ for instruction and $\mathcal{S}{\pi}^{\text{res.}}(X_{i},Y_{i})$ for instruction-response pair based on three key metrics $\pi$ from AlignBench [37]: Clarity, Completeness, and Factuality. The final quality score $\mathcal{S}(X_{i},Y_{i})$ is obtained by averaging these scores. These samples with scores below a similar dynamic threshold are identified as low-quality, forming the seed dataset $\mathcal{D}^{\text{low}}$.
@@
In each iteration, after the three data selection and optimization processes described above, the optimized dataset is then fed back for the next round of model training.
@@ 4 Experiment
4.1 Settings
Data Optimization Configurations.
We conduct optimization on the Alpaca [49] and WizardLM [55] datasets. For a fair comparison, we also include a rewritten version of Alpaca, where responses are generated by GPT-4o-mini, in our optimization process. Each dataset undergoes three iterations of optimization. Demonstrating that our method does not require a powerful external model, we synthesize data using DataDreamer [42] with GPT-4o-mini, setting both temperature and top_p to 1.0 to ensure diversity. A detailed analysis of the computational cost is provided in Appendix 7, and the effects of the number of neighbors and iteration counts are discussed in Appendix 8.
Training and Evaluation Settings.
We fine-tune LLaMA-3.1-8B [12] and Mistral-7B-v0.3 [20] using LLaMA-Factory [63] with the specific hyperparameters detailed in Appendix 9.5. For each iteration of Middo’s optimization, the base model is fine-tuned for one epoch on the dataset optimized in that specific iteration to mitigate the risk of overfitting to the data [26, 34]. Evaluation is conducted using OpenCompass [9], with vLLM [24] for acceleration. To validate the effectiveness and generalization capabilities of our approach, we assess model capabilities in general knowledge using IFEval [67] and MMLU [17]; mathematical problem-solving on GSM8K [7] and MATH [16]; code generation on HumanEval [6] and MBPP [2]; and commonsense reasoning on Hellaswag [59] and GPQA [45].
4.2 Main Results
@@
Further Validation on 4o-mini Rewritten Data.
Steady improvements observed on the 4o-mini rewritten Alpaca dataset—averaging a $2.2\%$ increase overall, with MMLU showing an impressive $11.87\%$ boost—demonstrate that these gains are not merely a result of using 4o-mini data. This illustrates that our framework intrinsically enhances data quality and model performance. Importantly, we achieve these improvements without resorting to stronger variants such as GPT-4o [18], reinforcing the robustness and general applicability of our method.
Initial Dataset Quality.
Our experiments reveal that higher-quality datasets require fewer modifications to reach optimal performance. On LLaMA-3.1-8B, for instance, while the Alpaca dataset achieves peak performance at third iteration, the 4o-mini rewritten Alpaca required only two iterations, and the Wizard dataset reaches its best performance in just one round.
Comparison with Other Works.
@@
Middo Alpaca 57.6kk 51.32 43.20 51.18 12.92 39.63 41.80 58.78 16.67 39.63
MiddOnly Alpaca 8.8kk 43.47 40.78 65.20 15.58 51.83 47.60 58.65 17.68 42.60
Middo Alpaca-4o-mini 63.1kk 44.69 47.96 57.62 18.50 52.44 45.40 57.37 19.70 42.96
MiddOnly Alpaca-4o-mini 24.9kk 41.50 45.66 60.80 20.06 46.34 48.00 55.01 24.75 42.77
Table 2: Results of Middo compared to other baseline methods. The best and second best results are highlighted in bold and underlined, respectively. Our method outperforms all baselines in the average score. *Note that 0.00 indicates that the method did not solve any examples. $\dagger$ Denotes training solely on Middo-generated data.
We compare Middo with both data selection (Alpaca-clean[46], Superfiltering [29], Long [61], AlpaGasus [5]) and data augmentation (Alpaca-GPT4 [43], I-SHEEP [34], WizardLM [55]) methods on the Alpaca dataset.
We use the optimal dataset obtained through Middo from Alpaca for comparison with other baselines. Additionally, to ensure a relatively fair comparison with data selection methods, we include a dataset that only uses the optimized data without incorporating any unoptimized samples, referred to as MiddOnly, to isolate the effect of the optimization process and make a direct comparison with data selection approaches.
Results in Table 2 show our method achieves the highest average score of $42.96$, outperforming all other approaches. Notably, even when using only the optimized subset MiddOnly Alpaca, our method delivers a robust average score of $42.60$. This demonstrates that iterative improvement is not primarily driven by data size, but rather by the effectiveness of our dynamic data selection and optimization process in identifying and generating data with high learning value for models.